Project

General

Profile

NDOS CommissionMtng 20101209

Attendance: Jon, Tom C., Leom, Kurt, Greg, Alena, Peter, Minerba, Andrew, Mark Messier, Chad, Athans, Art, Nate, Mayly

Agenda:
  1. DAQ/Electronics Update & Plans
  2. DCS Status & Plans
  3. Shifters' report
  4. AOB

DAQ:

  • Peter:
    • Currently testing fix to DCM firmware (improvement to data corruption).
      • DCM's FPGA reset is affecting FEB program.
      • A race condition caused imcomplete clearing of the FEB
      • Fixed now, but channel 3 in every DCM needs to be removed; this should be fixed in the next firmware release
  • Trigger timing:
    • Still don't know correct offset, but the plan is to determine the offset by looking at data offline
    • We will continue to run with no trigger offset but with a trigger window of 500 us (starting w/ run 10724). Current threshold is 500 ADC
  • Chiller: at the time, the chiller was being worked on (valve replacement).
  • The plan:
    • overnight: normal runs
      - daytime: DSO mode studies, DCM firmware development/testing

DCS:

  • Chad:
    • now able to connect to the dB in Labview and can read data from
      tables
    • working on getting (audible) alarms in the CR
    • Athans: has written test sender/receiver applications for DCM DCS monitor.

Shifter report:

nothing special to report.